# SKADA-Bench: Benchmarking Unsupervised Domain Adaptation Methods with Realistic Validation On Diverse Modalities

### v3

@@ Abstract
Unsupervised Domain Adaptation (DA) consists of adapting a model trained on a labeled source domain to perform well on an unlabeled target domain with some data distribution shift. While many methods have been proposed in the literature, fair and realistic evaluation remains an open question, particularly due to methodological difficulties in selecting hyperparameters in the unsupervised setting. With SKADA-Bench, we propose a framework to evaluate DA methods on diverse modalities, beyond computer vision task that have been largely explored in the literature. We present a complete and fair evaluation of existing shallow algorithms, including reweighting, mapping, and subspace alignment. Realistic hyperparameter selection is performed with nested cross-validation and various unsupervised model selection scores, on both simulated datasets with controlled shifts and real-world datasets across diverse modalities, such as images, text, biomedical, and tabular data. Our benchmark highlights the importance of realistic validation and provides practical guidance for real-life applications, with key insights into the choice and impact of model selection approaches. SKADA-Bench is open-source, reproducible, and can be easily extended with novel DA methods, datasets, and model selection criteria without requiring re-evaluating competitors. SKADA-Bench is available on Github at https://github.com/scikit-adaptation/skada-bench.
@@ 1 Introduction
Figure 1: Illustration of the different data shifts and assumptions studied in the DA literature and used in the simulated datasets used in the numerical experiments. The colors indicate the classes, with solid colors representing target and transparent colors representing source in the target data plot.
@@
Contributions.  In the following, we propose SKADA-Bench, an ambitious and fully reproducible benchmark with the following features: 1. A set of 4 simulated and 8 real-life datasets with different modalities (computer vision, NLP, tabular, biomedical) totaling 51 realistic shift scenarios, 2. A wide range of 20 Shallow DA methods designed to handle different types of shifts, 3. An evaluation of 3 deep DA methods on 4 real-world datasets from the computer vision and biomedical modalities, 4. A realistic model selection procedure using 5 different unsupervised scorers with nested cross-validation for hyperparameter selection, 5. An open-source implementation and publicly available datasets, easy to extend for new DA methods and datasets without the need to re-run the whole experiment.
@@ 3.1 Nested cross-validation loop and implementation
Python implementation. The benchmark code is available on GitHub upon publication of the paper.***Our code is available in supplementary materials. Our benchmark is implemented following the benchopt framework (Moreau et al., 2022), which provides standardized ways of organizing and running benchmarks for ML in Python. This framework facilitates reproducing the benchmark’s results, with tools to install the dependencies, run the methods in parallel, or cache the results to prevent redundant computations. It also makes it easy to extend the benchmark with additional datasets and methods, enabling it to evolve to account for the advances in the field. In the supplementary materials, we provide examples demonstrating how to add DA methods or datasets to the benchmark. Using this framework, we aim to make SKADA-Bench a reference benchmark to evaluate new DA methods in realistic scenarios with valid performance estimations.
@@ 3.2 Compared DA methods
Deep DA methods. These methods aim to reduce the divergence between the source and target data distributions within the learned feature space while simultaneously learning a classifier on source data. The training loss consists in a traditional supervised loss on labeled source data and a second term measuring the discrepancy between the source and target distributions. The methods implemented in the Deep DA Benchmark use different discrepancies, such as covariance distance (Sun & Saenko, 2016) for DeepCORAL, adversarial loss (Ganin et al., 2016a) for DANN and optimal transport distance (Damodaran et al., 2018a) for DeepJDOT. Note that these approaches are not part of the main shallow DA benchmark but have been added to provide an interesting comparison of DA performances between shallow and Deep methods on computer vision and biomedical data.
@@ 3.3 Compared datasets
Before using shallow DA methods, the datasets are preprocessed with feature extraction to ensure reasonable performance when trained on each domain. For example, images are embedded using Deep pre-trained models followed by a PCA (except MNIST/USPS where only PCA is used), and textual data is embedded using Large Language Models (LLM) (Reimers & Gurevych, 2019; Xiao et al., 2023a) before applying a PCA. The tabular data are one-hot encoded to transform categorical data into numerical data. The biosignals from Brain-Computer Interface (BCI) data are embedded using the state-of-the-art tangent space representation proposed in Barachant et al. (2012). For Deep DA methods, only 4 datasets are used: Office31, Office Home, MNIST/USPS and BCI. Since these methods focus on learning feature representations, the data are used in their raw form. The datasets are split into pairs of source and target domains totaling $51$ adaptation tasks in the benchmark. More details about the datasets and pre-processing are available in Appendix C.
@@ 4.1 Performance of the DA methods
Results table.
First, we report the realistic performances of the different methods when using their selected scorer on the different datasets in Table 2. The cells showcasing a significant change in performance with the Wilcoxon test are highlighted with colors. Blue indicates an increase in performance, while red indicates a loss. The intensity of the color corresponds to the magnitude of the gain or loss - the darker the shade, the larger the positive or negative change. Cells with a NA values indicate that the method was not applicable to the dataset (DASVM is limited to binary classification) or that the method has reached a timeout. We also report the best scorer and the average rank of the methods for all real datasets. In addition to Table 2 providing realistic performance estimations with the best realistic scorer, we also report in Table 20 (Appendix E) the results when using the non-realistic supervised scorer.
@@
Mapping CORAL 0.74 0.7 0.76 0.18 0.65 0.57 0.62 0.73 0.7 0.72 0.92 0.62 CircV 5.08
@@
Real data with unknown shift. The performance of reweighting methods is often close to Train Src baseline on real datasets. This result can be be due to the violations of the same-support assumption, which is crucial for reweighting to work effectively (Segovia-Martín et al., 2023) which is likely true for the three CV datasets. In this case, hyperparameter tuning frequently select configurations leading to near-uniform weighting, which explain the close performance to Train Src.
The performance of mapping methods is dataset-dependent, potentially due to the number of classes and presence of target shift. Mapping methods excel on MNIST/USPS and 20NewsGroup which respectively contain 10 and 2 classes, but failing on Office31 and OfficeHome with 31 and 60 classes. Additionally, while mapping performs well on the NLP dataset 20NewsGroup, it results in negative transfer on Amazon Reviews which has target shifts.
It is notable that simple transformations are the best in average across all modalities. Indeed most methods that significantly outperform Train Src in ranking average across all modalities are LinOT, CORAL, JPCA, and SA, which all rely on linear transformations such as scaling, linear projection or rotations. These methods are robust across datasets and modalities, offering effective alignment with minimal risk of negative DA.
@@
Selected scorer per DA method. We observe that the best scorer differs across methods, Circular Validation has been selected 10 out of 20 times as the best scorer, followed by Importance Weighting 4 out of 20 times. Table 20 in the supplementary material provides the non-realistic accuracy results with the supervised scorer. It is worth noting that the supervised scorer generally outperforms the unsupervised ones, and several methods significantly outperform Train Src in each dataset. It is crucial to choose the model realistically to avoid producing overly optimistic results, as many data analysis papers have done (see Table 4).
@@ 4.2 Study of validation scorers
Furthermore, we provide in Figures 8 and 9, from Appendix E, several visualizations that illustrate the relationship between the accuracy achieved when using a supervised scorer and the accuracy obtained when using different unsupervised scorers. We also visualize in Figure 7 the drop in performance when using the best-unsupervised scorer instead of the supervised scorer. Interestingly some methods such as KMM, EntOT, and ClassRegOT can lose up to 10% accuracy when using realistic scorers, which might come from their higher number of parameters or their sensitivity to them.
Our results thus show that most scorers have poor results when evaluated on many datasets. Of the five methods under consideration, only two achieve satisfactory performance, although incurring large variance in their results. This shows that proper hyperparameter selection is still an open question, that needs attention from the research comunity to guide practitioners toward real life applications of unsupervised DA technics.
@@ 4.3 Deep DA methods
The results are available in Table 3 with a comparison to the best performing shallow method from Table 2. One of the most notable and expected difference is on MNIST/USPS. Shallow methods struggle to achieve good performances, even on Train Tgt, as they rely on PCA for feature extraction. Deep methods, on the other hand, use CNNs, leading to large accuracy gains on train on Src and Tgt but also on Deep DA methods. However, it is important to note that while DeepJDOT, DANN and MCC improve performance on all datasets, they remain far from the train on Tgt accuracies, partly due to the difficulty in tuning their parameters (see Appendix E.3 with the supervised scorer). The superior performance of Deep DA methods on CV datasets can be attributed to the relationship between classification in the DA subspace and the disentanglement of semantic (discriminant) content from style (domain shift) (Gonzalez-Garcia et al., 2018; Gabbay et al., 2021). Numerous studies have demonstrated that semantic embeddings can be effectively recovered, supporting the assumption that a (nonlinear) subspace shift is reasonable for CV tasks. However, for the BCI dataset, where the amount of data is limited, the performances of Deep DA methods are inferior to some other shallow methods (i.e., LinOT for example). Finally, a method like DANN, which is often considered as a baseline in the community, has been shown to be hard to validate and requires setup that can be difficult to determine across different settings. These results emphasize, that while Deep invariant DA methods can be effective, they do not consistently yield good results across modalities, whereas shallow DA methods can achieve similar or superior performances with less effort and fewer computational resources in low data regimes.
@@ Acknowledgments
This work was supported by grant ANR-22-PESN-0012 to AC under the France 2030 program, ANR-20-CHIA-0016 and ANR-20-IADJ-0002 to AG, and ANR-23-ERCC-0006 to YL, TG, and RF, all managed by the Agence Nationale de la Recherche (ANR). Additionally, this project received funding from the European Union’s Horizon Europe research and innovation program under grant agreement 101120237 (ELIAS).
@@ Appendix A Appendix
The entire code and results of SKADA-Bench will be open-sourced and available online. The implementation of the DA methods and scorers is provided along with access to the simulated and real-world datasets. All the performance tables and figures can be reproduced effortlessly, and guidelines with minimal working examples are given to add new DA methods and datasets.
@@ Appendix B Model selection in Domain Adaptation
Reweighting Density Reweight  (Sugiyama & Müller, 2005) None Bandwidth fixed by Silverman method
Discriminative Reweight  (Shimodaira, 2000) NA No hyperparameters
Gaussian Reweight  (Shimodaira, 2000) None Not specified in  (Shimodaira, 2000)
KLIEP  (Sugiyama et al., 2007b) Integrated CV Likelihood CV (Sugiyama et al., 2007b) on target
KMM  (Huang et al., 2006) None Fixed data-dependent hyperparameters
NN Reweight  (Loog, 2012) None Number of neighbors fixed to one
MMDTarS  (Zhang et al., 2013) CV Not specified if done on source or target
Mapping Coral  (Sun et al., 2017) NA No hyperparameters
OT mapping  (Courty et al., 2017b) CV target/CircCV Unclear in the text
Lin. OT mapping  (Flamary et al., 2020) NA No hyperparameters
MMD-LS  (Zhang et al., 2013) CV Not specified if done on source or target
Subsp. SA  (Fernando et al., 2013) 2-fold CV on source -
TCA  (Pan et al., 2011) Validation on target Target subset used to tune parameters
TSL  (Si et al., 2010) None Not specified in (Si et al., 2010)
Other JDOT  (Courty et al., 2017a) Reverse CV  (Zhong et al., 2010) -
OT label prop  (Solomon et al., 2014) NA No hyperparameters
DASVM  (Bruzzone & Marconcini, 2010a) Circular Validation  (Bruzzone & Marconcini, 2010a) -
@@ E.1 Results per datasets
These detailed tables where cell in green denote a gain wrt Train Src (average outside of standard deviation of Train Src) better illustrate the challenges of domain adaptation (DA) methods. They show that not all shifts are equivalent within a given dataset. For example, Table 12 reveals that only 3 shifts in the AmazonReview dataset present a DA problem (defined as a $>3\%$ difference in accuracy between Train Src and Train Tgt). While for the other shifts, we achieve similar performance whether we train on source or target data. Additionally, some specific shifts present a DA problem that no method can successfully address. This can be seen in the dsl → amz shift in the Office31 dataset, as shown in Table 7. Finally, some DA methods perform consistently across all shifts within a dataset, as demonstrated by the results for the 20Newsgroup dataset in Table 11.
@@
[…]
@@
[…]

### v4

@@ Abstract
Unsupervised Domain Adaptation (DA) consists of adapting a model trained on a labeled source domain to perform well on an unlabeled target domain with some data distribution shift. While many methods have been proposed in the literature, fair and realistic evaluation remains an open question, particularly due to methodological difficulties in selecting hyperparameters in the unsupervised setting. With SKADA-Bench, we propose a framework to evaluate DA methods on diverse modalities, beyond computer vision task that have been largely explored in the literature. We present a complete and fair evaluation of existing shallow algorithms, including reweighting, mapping, and subspace alignment. Realistic hyperparameter selection is performed with nested cross-validation and various unsupervised model selection scores, on both simulated datasets with controlled shifts and real-world datasets across diverse modalities, such as images, text, biomedical, and tabular data. Our benchmark highlights the importance of realistic validation and provides practical guidance for real-life applications, with key insights into the choice and impact of model selection approaches. SKADA-Bench is open-source, reproducible, and can be easily extended with novel DA methods, datasets, and model selection criteria without requiring re-evaluating competitors. The code is available at https://github.com/scikit-adaptation/skada-bench.
@@ 1 Introduction
Figure 1: Illustration of different type of distribution shift between source and target domains: covariate shift, target shift, conditional shift, and subspace shift (Mathematial details are available in Section 2.1). Points represent data samples, with colors indicating different classes. These synthetic datasets are used to evaluate model performance under controlled shift scenarios in the experiement part.
@@
Contributions.  In the following, we propose SKADA-Bench, an ambitious and fully reproducible benchmark with the following features: 1. A set of 4 simulated and 8 real-life datasets with different modalities (computer vision, NLP, tabular, biomedical) totaling 51 realistic shift scenarios, 2. A wide range of 20 Shallow DA methods designed to handle different types of shifts, 3. An evaluation of 7 deep DA methods on 4 real-world datasets from the computer vision and biomedical modalities, 4. A realistic model selection procedure using 5 different unsupervised scorers with nested cross-validation for hyperparameter selection, 5. An open-source implementation and publicly available datasets, easy to extend for new DA methods and datasets without the need to re-run the whole experiment.
@@ 3.1 Nested cross-validation loop and implementation
Python implementation. The codebase is available at https://github.com/scikit-adaptation/skada-bench. Our benchmark is implemented following the benchopt framework (Moreau et al., 2022), which provides standardized ways of organizing and running benchmarks for ML in Python. This framework facilitates reproducing the benchmark’s results, with tools to install the dependencies, run the methods in parallel, or cache the results to prevent redundant computations. It also makes it easy to extend the benchmark with additional datasets and methods, enabling it to evolve to account for the advances in the field. In the supplementary materials, we provide examples demonstrating how to add DA methods or datasets to the benchmark. Using this framework, we aim to make SKADA-Bench a reference benchmark to evaluate new DA methods in realistic scenarios with valid performance estimations.
@@ 3.2 Compared DA methods
Deep DA methods. These methods aim to reduce the divergence between the source and target data distributions within the learned feature space while simultaneously learning a classifier on source data. The training loss consists in a traditional supervised loss on labeled source data and a second term measuring the discrepancy between the source and target distributions. The methods implemented in the Deep DA Benchmark use different discrepancies, such as covariance distance (Sun & Saenko, 2016) for DeepCORAL, adversarial loss (Ganin et al., 2016a; Zhang et al., 2019) for DANN or MDD, MMD distance (Long et al., 2015a) for DAN, optimal transport distance (Damodaran et al., 2018a) for DeepJDOT, class confusion (Jin et al., 2020) for MCC, and graph spectral alignment (Xiao et al., 2023b) for SPA. Note that these approaches are not part of the main shallow DA benchmark but have been added to provide an interesting comparison of DA performances between shallow and Deep methods on computer vision and biomedical data.
@@ 3.3 Compared datasets
Data are preprocessed to extract relevant features, while keeping computational costs reasonable. Images are embedded using deep pre-trained models (except MNIST/USPS where the images are vectorized), and textual data is embedded using Large Language Models (LLM) (Reimers & Gurevych, 2019; Xiao et al., 2023a). The tabular data are one-hot encoded to transform categorical data into numerical data. The biosignals from Brain-Computer Interface (BCI) data are embedded using the state-of-the-art tangent space representation proposed in Barachant et al. (2012). For images and text, we apply a PCA to reduce the dimensionality of the embeddings to a reasonable dimension to avoid big computational costs. For Deep DA methods, only 4 datasets are used: Office31, Office Home, MNIST/USPS and BCI. Since these methods focus on learning feature representations, the data are used in their raw form. The datasets are split into pairs of source and target domains totaling $51$ adaptation tasks in the benchmark. More details about the datasets and pre-processing are available in Appendix C.
@@ 4.1 Performance of the DA methods
Results table. First, we report the realistic performances of the different methods when using their selected scorer on the different datasets in Table 2. The cells showcasing a significant change in performance with the Wilcoxon test are highlighted with colors. Blue indicates an increase in performance, while red indicates a loss. The intensity of the color corresponds to the magnitude of the gain or loss - the darker the shade, the larger the positive or negative change. Cells with a NA values indicate that the method was not applicable to the dataset (DASVM is limited to binary classification) or that the method has reached a timeout. We also report the best scorer and the average rank of the methods for all real datasets. In addition to Table 2 providing realistic performance estimations with the best realistic scorer, we also report in Table 20 (Appendix E) the results when using the non-realistic supervised scorer.
@@
Mapping
CORAL 0.74 0.7 0.76 0.18 0.65 0.57 0.62 0.73 0.7 0.72 0.92 0.62 CircV 5.08
@@
Real data with unknown shift. The performance of reweighting methods is often close to Train Src baseline on real datasets. For example, NN RW reaches only a 54% accuracy on MNIST/USPS compared to 54% for Train Src. This result can be be due to the violations of the same-support assumption, which is crucial for reweighting to work effectively (Segovia-Martín et al., 2023), and which is likely true for the three CV datasets. In this case, hyperparameter tuning frequently select configurations leading to near-uniform weighting, which explain the close performance to Train Src.
The performance of mapping methods is dataset-dependent, potentially due to the number of classes and presence of target shift. Mapping methods excel on MNIST/USPS and 20NewsGroup which respectively contain 10 and 2 classes, but fail on Office31 and OfficeHome with 31 and 60 classes. For example, ClassRegOT achieves 62% on MNIST/USPS and 97% on 20NewsGroup, but only 53% on OfficeHome. Additionally, while mapping performs well on the NLP dataset 20NewsGroup, it results in negative transfer on Amazon Reviews which has target shifts.
It is notable that simple transformations are the best in average across all modalities. Indeed, most methods that significantly outperform Train Src in ranking average across all modalities are LinOT, CORAL, JPCA, and SA, which all rely on linear transformations such as scaling, linear projection, or rotations. These methods are robust across datasets and modalities, offering effective alignment with minimal risk of negative DA. For instance, LinOT consistently ranks among the top 5 methods and achieves substantial gains over Train Src: +10 points on MNIST/USPS (64% vs. 54%), +23 points on 20NewsGroup (82% vs. 59%), and +6 points on BCI (61% vs. 55%).
Computational cost. While full computational results are reported in Appendix E.7 and Figure 11, we briefly highlight that several top-performing methods—such as LinOT, CORAL, and SA—also exhibit some of the lowest training and testing times, averaging under 20 seconds per task. This reinforces their appeal in practice: they combine competitive performance with high efficiency, making them especially suitable for settings with limited computational resources or many adaptation tasks.
@@
Selected scorer per DA method. We observe that the best scorer differs across methods, Circular Validation has been selected 10 out of 20 times as the best scorer, followed by Importance Weighting 4 out of 20 times. Table 20 in the supplementary material provides the non-realistic accuracy results with the supervised scorer. It is worth noting that the supervised scorer generally outperforms the unsupervised ones. For example, EntOT achieves +5 points with the supervised scorer versus with the CircV scorer (88% vs 83%) on the 20NewsGroups dataset. It is crucial to choose the model realistically to avoid producing overly optimistic results, as many data analysis papers have done (see Table 4).
@@ 4.2 Study of validation scorers
Furthermore, we provide in Figures 9 and 10, from Appendix E, several visualizations that illustrate the relationship between the accuracy achieved when using a supervised scorer and the accuracy obtained when using different unsupervised scorers. We also visualize in Figure 8 the drop in performance when using the best-unsupervised scorer instead of the supervised scorer. Interestingly some methods such as KMM, EntOT, and ClassRegOT can lose up to 10% accuracy when using realistic scorers, which might come from their higher number of parameters or their sensitivity to them.
Our results thus show that most scorers have poor results when evaluated on many datasets. Of the six methods under consideration, only two achieve satisfactory performance, although incurring large variance in their results. This shows that proper hyperparameter selection is still an open question, that needs attention from the research comunity to guide practitioners toward real life applications of unsupervised DA technics.
@@ 4.3 Deep DA methods
Results table. The results are available in Table 3 with a comparison to the best performing shallow method from Table 2. One of the most notable and expected difference is on MNIST/USPS. Shallow methods struggle to achieve good performances, even on Train Tgt, as they rely on PCA for feature extraction. Deep methods, on the other hand, use CNNs, leading to large accuracy gains on train on Src and Tgt but also on Deep DA methods. However, it is important to note that while DeepJDOT, DANN and MCC improve performance on all datasets, they remain far from the train on Tgt accuracies, partly due to the difficulty in tuning their parameters (see Appendix E.3 with the supervised scorer). The superior performance of Deep DA methods on CV datasets can be attributed to the relationship between classification in the DA subspace and the disentanglement of semantic (discriminant) content from style (domain shift) (Gonzalez-Garcia et al., 2018; Gabbay et al., 2021). Numerous studies have demonstrated that semantic embeddings can be effectively recovered, supporting the assumption that a (nonlinear) subspace shift is reasonable for CV tasks. However, for the BCI dataset, where the amount of data is limited, the performances of Deep DA methods are inferior to some other shallow methods (i.e., LinOT for example). Finally, a method like DANN, which is often considered as a baseline in the community, has been shown to be hard to validate and requires setup that can be difficult to determine across different settings. These results emphasize, that while Deep invariant DA methods can be effective, they do not consistently yield good results across modalities, whereas shallow DA methods can achieve similar or superior performances with less effort and fewer computational resources in low data regimes.
Limitations and future work.
The evaluation of deep DA methods in this benchmark is limited to a single outer split to ensure computational feasibility. While practical, this setting may introduce variance in performance estimates, especially on small datasets. Moreover, although all deep methods completed within the allocated time, the 4-hour timeout may restrict the exploration of more complex architectures or longer training procedures on larger datasets.
We view these constraints as a limitation of the current benchmark and an opportunity for future work. A comprehensive evaluation of deep DA methods would benefit from larger datasets ($10^{4}$ samples) with more stable training dynamics, and lower-variance evaluation using a single train/validation/test split. Extending DA-Bench to support such large-scale evaluations represents a promising direction for improving the practical assessment of deep DA approaches.
@@ Acknowledgments
This work benefited from state aid managed by the Agence Nationale de la Recherche under the France 2030 programme, reference ANR-23-IACL-0005, ANR-22-PESN-0012 and ANR-20-CHIA-0016. This research was also supported in part by the French National Research Agency (ANR) through the MATTER project (ANR-23-ERCC-0006-01) and the BenchArk project (ANR-24-IAS2-0003). It received funding from the Fondation de l’École polytechnique. Additionally, this project received funding from the European Union’s Horizon Europe research and innovation program under grant agreement 101120237 (ELIAS).
@@ Appendix A Appendix
The entire code and results of SKADA-Bench are open-sourced at https://github.com/scikit-adaptation/skada-bench. The implementation of the DA methods and scorers is provided along with access to the simulated and real-world datasets. All the performance tables and figures can be reproduced effortlessly, and guidelines with minimal working examples are given to add new DA methods and datasets.
@@ Appendix B Model selection in Domain Adaptation
Reweighting
Density Reweight
 (Sugiyama & Müller, 2005)
None
Bandwidth fixed
by Silverman method
Discriminative Reweight
 (Shimodaira, 2000)
NA No hyperparameters
Gaussian Reweight
 (Shimodaira, 2000)
None
Not specified in
 (Shimodaira, 2000)
KLIEP
 (Sugiyama et al., 2007b)
Integrated CV
Likelihood CV
(Sugiyama et al., 2007b)
on target
KMM
 (Huang et al., 2006)
None
Fixed data-dependent
hyperparameters
NN Reweight
 (Loog, 2012)
None
Number of neighbors
fixed to one
MMDTarS
 (Zhang et al., 2013)
CV
Not specified if
done on source or target
Mapping
Coral
 (Sun et al., 2017)
NA No hyperparameters
OT mapping
 (Courty et al., 2017b)
CV target/CircCV Unclear in the text
Lin. OT mapping
 (Flamary et al., 2020)
NA No hyperparameters
MMD-LS
 (Zhang et al., 2013)
CV
Not specified if
done on source or target
Subsp.
SA
 (Fernando et al., 2013)
2-fold CV on source -
TCA
 (Pan et al., 2011)
Validation on target
Target subset used
to tune parameters
TSL
 (Si et al., 2010)
None
Not specified
in (Si et al., 2010)
Other
JDOT
 (Courty et al., 2017a)
Reverse CV
 (Zhong et al., 2010)
OT label prop
 (Solomon et al., 2014)
NA No hyperparameters
DASVM
 (Bruzzone & Marconcini, 2010a)
Circular Validation
 (Bruzzone & Marconcini, 2010a)
@@ E.1 Results per datasets
These detailed tables where cell in green denote a gain wrt Train Src (average outside of standard deviation of Train Src) better illustrate the challenges of domain adaptation (DA) methods. They show that not all shifts are equivalent within a given dataset. For example, Table 12 reveals that only 3 shifts in the AmazonReview dataset present a DA problem (defined as a $3\%$ difference in accuracy between Train Src and Train Tgt). While for the other shifts, we achieve similar performance whether we train on source or target data. Additionally, some specific shifts present a DA problem that no method can successfully address. This can be seen in the dsl → amz shift in the Office31 dataset, as shown in Table 7. Finally, some DA methods perform consistently across all shifts within a dataset, as demonstrated by the results for the 20Newsgroup dataset in Table 11.
@@
$\rightarrowdsl amz\rightarrowweb$ $\rightarrowamz dsl\rightarrowweb$ $\rightarrowamz web\rightarrowdsl$ Mean Rank
@@
$\rightarrowsci rec\rightarrowtalk$ $\rightarrowrec sci\rightarrowtalk$ $\rightarrowrec talk\rightarrowsci$ Mean Rank